# CaloHadronic: a diffusion model for the generation of hadronic showers

### v1

@@ 1 Introduction
As experiments in high energy physics push the boundaries of luminosity resulting in ever increasing event rates, the computational demand of high-precision Monte Carlo (MC) simulations is growing to the point where it will soon surpass available computational resources [4]. Generative models offer a promising solution to this challenge, potentially reducing the immense computational load required for these simulations. This has led to substantial research into the development of machine-learning architectures tailored for more efficient and accurate detector simulation [5, 6]. Examples include generative adversarial networks (GANs) [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], variational autoencoders (VAEs) and their variants [19, 20, 21, 18, 22, 23], normalizing flows and various types of diffusion models  [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 23, 39, 40, 41, 42, 43, 44], as well as generative pre-trained transformer (GPT) style models [45]. The combination of a diffusion model with a transformer architecture, known as diffusion transformers [46, 47], has been used in high-energy physics for jet generation [48, 49, 50, 51, 44].
The majority of these studies have focused on simulating electromagnetic showers. Only a few of them (e.g. [21]) have attempted to model hadronic showers, which exhibit significantly more complex and varied phenomenology with respect to electromagnetic ones. Additionally, studies have focused either on simulating showers in the electromagnetic calorimeter (ECal) or in the hadronic calorimeter (HCal) with high precision. In practice, however, a hadronic particle might already start interacting in the electromagnetic calorimeter and then continue showering in the hadronic calorimeter (and perhaps even beyond). Therefore, for realistic applications of these models, an efficient and comprehensive simulation of the entire shower is necessary. This work represents the first approach to provide a unified framework for modeling hadronic showers across both the electromagnetic and hadronic components of a highly granular calorimeter system.
As electromagnetic and hadronic calorimeters typically have different cell sizes [52] — leading to different spatial resolutions or, put differently, to a change in diameter of the shower when measured in cells (e.g. [53]) — and are built from different materials, as [54], correctly learning the behavior across the two detector systems is not trivial. Together with the much larger spatial extent and the much larger statistical fluctuations of hadronic showers compared to electromagnetic ones, this represents a major challenge for approaches based on regular grids.
@@ 2.1 Data Preprocessing
Following the approach in Ref. [30], all Geant4 steps are extracted from the sensitive layers of both calorimeters, yielding 2D point clouds for every one of the 78 layers. Subsequently, each layer is projected onto a virtual grid with roughly 9 times higher granularity than the original ECal and HCal: 1.7x1.7 mm2 granularity is used for the ECal, while a coarser granularity of 10x10 mm2 is used for the HCal. Consequently, the total number of grid points per layer is $280,900$ for the ECal and $8,100$ for the HCal, given a cut along the transverse plane (square cut in the $x$-$z$ plane) from -450 mm to 450 mm. In a similar fashion to Ref. [30] all layers are stacked along the $y$-axis, with the $y$-coordinate assigned according to the layer number. A random offset uniformly sampled from the interval $[-0.5,0.5]$ is applied to each point’s $y$-position, dequantizing the longitudinal shower profile.
@@
After model inference, the individual energy depositions produced by the model are positioned into the detector geometry using the DDML444 https://github.com/key4hep/DDML/ library [61, 62], which uses the fast simulation hooks present in Geant4 and DD4hep. This enables an efficient placement and scoring of hits in the sensitive elements of the detector.
@@ 3.2 ECal and HCal blocks
An illustration of ECal edm-diffusion is shown on the left of figure 1. For ECal edm-diffusion, the conditional features are the incident energy and the number of points per layer for each of the 30 layers of the electromagnetic calorimeter of ILD. In their embedding, both of them are concatenated and then passed through two fully connected layers. The time feature of the diffusion model also has an embedding, consisting of a Fourier feature mapping and then two fully connected layers. The Fourier mapping of the time variable is taken from [56] and is similar to the previously mentioned Fourier Layer. The 4D input, consisting of the three coordinates ($x$,$y$ and $z$) and the energy ($e$) of the input shower, goes into custom defined embedding layers based on the Fourier Layer and then two fully connected layers as well. Time, conditioning and input embeddings, with a dimension of 128, are concatenated along the points axis and fed into three layers of transformer encoder [46]. After removing the conditioning points, a projection with three fully connected layers is applied to map to the desired output dimension of four features: $x$, $y$, $z$, $energy$.
Figure 1: Illustration of the structure of the ECal edm-diffusion (left) and HCal edm-diffusion (right). $time$ is the time of the diffusion model used to add noise to the data and $cond.$ are the incident energy and the number of points per layer. The electromagnetic calorimeter (ECal) consists of 30 layers, and the hadronic calorimeter (HCal) consists of 48 layers. input (ECal or HCal) is the input shower to be learned during training.
@@ 4 Results
[…]
@@
Figure 3 shows that the tracks resulting from hadronic interactions occurring in $\pi^{+}$ showers, are well reproduced by CaloHadronic. These was possible due to the self-attention mechanism that took care of the interactions among points. The shower core is well defined and the tracks are visible. Notably, the CaloHadronic model better reproduces fine-grained, track-like structures in the HCal, which are especially evident when compared to the Geant4 reference. Noise manifests as low-energy deposits spread outside the core of the shower and can arise from detector effects such as electronics noise, cross-talk, or background activity. More generated $\pi^{+}$ showers can be found in appendix D.
@@ 4.1 Physics Performance
The per-cell energy distribution (figure 4, left) contains the energy of the cells of all showers in the test dataset. In this distribution there are two peaks. The first one at $\sim$0.2 MeV corresponds to the most probable energy deposition of a minimum ionizing particle (MIP) in the silicon sensors of the ECal while the second one at $\sim$1 MeV corresponds to the MIP of the HCal. The model describes this distribution reasonably well. The largest discrepancy appears for higher cell energies where the statistics are lower.
The longitudinal shower profile (figure 4, center) describes how much energy is deposited on average in each of the 78 calorimeter layers. The first 30 layers are part of the ECal, while the last 48 compose the HCal. The model describes the distribution well. It can be observed that there are small bumps in the ECal layers, where the Geant4 simulation differs between even and odd layers. This alternating structure of higher and lower energy depositions per layer arises from the fact that for technical reasons, pairs of silicon sensors are placed on either side of one tungsten absorber layer facing opposite directions. These slabs are then installed into a tungsten structure containing every other absorber layer. The different amount of nonsensitive material results in the observed pair-wise difference in the mean energy between consecutive layers. CaloHadronic has some difficulties in fully modeling this difference.
The radial shower profile (figure 4, right) describes the average radial energy distribution around the central shower axis (in the $y$-direction) in both ECal and HCal. CaloHadronic has a tendency to generate slightly narrower showers than Geant4.
@@
In figure 7 the energy sum (total visible energy) for 5k showers at three incident energies for ECal, HCal and both calorimeters combined is shown.
@@ 4.2 Correlation Studies
Figure 11 shows two 2D histograms of the total visible energy deposited in the ECal against the total visible energy deposited in the HCal for incident energies of 15 GeV, 50 GeV and 85 GeV, respectively. The 1D histogram on the right side gives a clearer view of the HCal total visible energy when the ECal total visible energy is below 38 MeV. Most of the pion showers lie in the region of lower energy depositions in the ECal and higher energy depositions in the HCal. This is due to the fact that a large fraction of the pions travel through the ECal before showering in the HCal, where the majority of their energy is deposited.
@@ 5.1 Methods
In order to be able to apply the standard reconstruction suite of ILD to showers produced by the model, it is necessary to be able to correctly place the energy deposits into the detector geometry. We therefore use the DDML library [61, 62] to combine the model output with a full simulation application. We fire a $\pi^{+}$ with an energy of $50$ GeV from a particle gun positioned at ($-5,0,-15$) mm [^6] This position is chosen to avoid the TPC cathode, which is positioned at $z=0$ in the global ILD coordinate system in the global ILD coordinate system, with a direction chosen to produce a perpendicular incidence at the calorimeter face. This takes into account the slight curvature of the track. The particles are not produced directly at the face of the calorimeter, as it is necessary for a track to be present in the tracking system for PandoraPFA to apply the correct sequence of algorithms for $\pi^{+}$ reconstruction. However, the energy of the pion used for conditioning of the model is extracted directly at the face of the calorimeter, to match the training scenario.
@@ 6 Conclusion
To turn generative models from proof-of-concept studies into useful surrogates to be used in production, they need to be able to handle the various complexities of realistic calorimeters. This work demonstrates for the first time how a point-cloud based generative model can be extended to simultaneously simulate the ECal and HCal parts of a hadronic shower, emphasizing the flexibility of this approach. Both individual shower properties, as well as correlations (also across ECal and HCal components), are described well. A classifier test shows that — as for the generation of partial showers — the generative fidelity needs to be further improved. This is especially visible in the fine grained track-like structures observed in the HCal. However the degree of agreement between generative output and ground-truth simulation is of a similar or slightly better quality when also considering down-stream simulation.
@@ 7 Acknowledgments
This research was supported in part by the Maxwell computational resources operated at Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany. This project has received funding from the European Union’s Horizon 2020 Research and Innovation programme under Grant Agreement No 101004761. We acknowledge support by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and via the KISS consortium (05D23GU4, 13D22CH5) funded by the German Federal Ministry of Research, Technology and Space (BMFTR) in the ErUM-Data action plan. P.M. has benefited from support by the CERN Strategic R&\&&D Programme on Technologies for Future Experiments [83].
@@ Appendix E CaloHadronic further histograms
Figure 20: Distribution of hit $x$ (left), $y$ (midel), and $z$ (right) positions calculated from $50000$ showers. All distributions are calculated with a uniform distribution of incident particle energies between 10 and 90 GeV. The error band corresponds to the statistical uncertainty in each bin.

### v2

@@ 1 Introduction
As experiments in high energy physics push the boundaries of luminosity resulting in ever increasing event rates, the computational demand of high-precision Monte Carlo (MC) simulations is growing to the point where it will soon surpass available computational resources [4]. Generative models offer a promising solution to this challenge, potentially reducing the immense computational load required for these simulations. This has led to substantial research into the development of machine-learning architectures tailored for more efficient and accurate detector simulation [8, 49]. Examples include generative adversarial networks (GANs) [69, 70, 29, 33, 66, 34, 10, 22, 42, 12, 38, 1], variational autoencoders (VAEs) and their variants [18, 17, 19, 1, 27, 31, 40], normalizing flows and various types of diffusion models  [82, 83, 84, 41, 85, 62, 16, 3, 63, 6, 24, 52, 51, 79, 50, 31, 90, 15, 35, 68, 36, 14], as well as generative pre-trained transformer (GPT) style models [13]. The combination of a diffusion model with a transformer architecture, known as diffusion transformers [89, 73], has been used in high-energy physics for jet generation [54, 61, 23, 64, 14].
The majority of these studies have focused on simulating electromagnetic showers, for a recent review see [39]. Only a few of them (e.g. [19]) have attempted to model hadronic showers, which exhibit significantly more complex and varied phenomenology with respect to electromagnetic ones. Additionally, studies have focused either on simulating showers in the electromagnetic calorimeter (ECal) or in the hadronic calorimeter (HCal) with high precision. In practice, however, a hadronic particle might already start interacting in the electromagnetic calorimeter and then continue showering in the hadronic calorimeter (and perhaps even beyond). Therefore, for realistic applications of these models, an efficient and comprehensive simulation of the entire shower is necessary. This work represents the first approach to provide a unified framework for modeling hadronic showers across both the electromagnetic and hadronic components of a highly granular calorimeter system.
As electromagnetic and hadronic calorimeters typically have different cell sizes [30] — leading to different spatial resolutions or, put differently, to a change in diameter of the shower when measured in cells (e.g. [74]) — and are built from different materials (e.g. [26]), correctly learning the behavior across the two detector systems is not trivial. Together with the much larger spatial extent and the much larger statistical fluctuations of hadronic showers compared to electromagnetic ones, this represents a major challenge for approaches based on regular grids.
@@ 2.1 Data Preprocessing
Following the approach in Ref. [16], all Geant4 steps are extracted from the sensitive layers of both calorimeters, yielding 2D point clouds for every one of the 78 layers. Subsequently, each layer is projected onto a virtual grid with roughly 9 times higher granularity than the original ECal and HCal: 1.7x1.7 mm2 granularity is used for the ECal, while a coarser granularity of 10x10 mm2 is used for the HCal. Consequently, the total number of grid points per layer is $280,900$ for the ECal and $8,100$ for the HCal, given a cut along the transverse plane (square cut in the $x$-$z$ plane) from -450 mm to 450 mm. In a similar fashion to Ref. [16] all layers are stacked along the $y$-axis, with the $y$-coordinate assigned according to the layer number. A random offset uniformly sampled from the interval [0.5,0.5][-0.5,0.5] is applied to each point’s $y$-position, dequantizing the longitudinal shower profile.
@@
After model inference, the individual energy depositions produced by the model are positioned into the detector geometry using the DDML [^4] https://github.com/key4hep/DDML/ library [59, 60], which uses the fast simulation hooks present in Geant4 and DD4hep. This enables an efficient placement and scoring of hits in the sensitive elements of the detector.
@@ 3.2 ECal and HCal blocks
An illustration of ECal edm-diffusion is shown on the left of figure 1. For ECal edm-diffusion, the conditional features are the incident energy and the number of points per layer for each of the 30 layers of the electromagnetic calorimeter of ILD. In their embedding, both of them are concatenated and then passed through two fully connected layers. The time feature of the diffusion model also has an embedding, consisting of a Fourier feature mapping and then two fully connected layers. The Fourier mapping of the time variable is taken from [46] and is similar to the previously mentioned Fourier Layer. The 4D input, consisting of the three coordinates ($x$,$y$ and $z$) and the energy ($e$) of the input shower, goes into custom defined embedding layers based on the Fourier Layer and then two fully connected layers as well. Time, conditioning and input embeddings, with a dimension of 128, are concatenated along the points axis and fed into three layers of transformer encoder [89]. After removing the conditioning points, a projection with three fully connected layers is applied to map to the desired output dimension of four features: $x$, $y$, $z$, energyenergy.
Figure 1: Illustration of the structure of the ECal edm-diffusion (left) and HCal edm-diffusion (right). timetime is the time of the diffusion model used to add noise to the data and $cond.$ are the incident energy and the number of points per layer. The electromagnetic calorimeter (ECal) consists of 30 layers, and the hadronic calorimeter (HCal) consists of 48 layers. input (ECal or HCal) is the input shower to be learned during training.
@@ 4 Results
The final point cloud showers are projected back to a regular grid in which the cell size is the same as that of the ILD calorimeters, hence 5x5 mm2 for the ECal and 30x30 mm2 for the HCal. For downstream analyses a cell energy cut at $\sim$$10^{-2}$ MeV is applied, since below this threshold the sensor response is indistinguishable from electronic noise. This cut was applied to all showers when calculating the shower observables and scores in this section. A first naive comparison of the showers generated by CaloHadronic with Geant4 simulated $\pi^{+}$ showers can be made from a visual inspection of whether the final projection of the showers resembles the Geant4 equivalent. In figure 3 the 3D view of a 50 GeV Geant4 shower (left) and a 50 GeV CaloHadronic shower (right) are shown.
@@
Figure 3 shows that the tracks resulting from hadronic interactions occurring in $\pi^{+}$ showers, are well reproduced by CaloHadronic. This was possible due to the self-attention mechanism that took care of the interactions among points. The shower core is well defined and the tracks are visible. Notably, the CaloHadronic model better reproduces fine-grained, track-like structures in the HCal, which are especially evident when compared to the Geant4 reference. Noise manifests as low-energy deposits spread outside the core of the shower and can arise from detector effects such as electronics noise, cross-talk, or background activity. More generated $\pi^{+}$ showers can be found in appendix D.
@@ 4.1 Physics Performance
The per-cell energy distribution (figure 4, left) contains the energy of the cells of all showers in the test dataset. In this distribution there are two peaks. The first one at $\sim$0.2 MeV corresponds to the most probable energy deposition of a minimum ionizing particle (MIP) in the silicon sensors of the ECal while the second one at $\sim$1 MeV corresponds to the MIP of the HCal. The model reproduces both peaks and the overall shape of the distribution reasonably well. The largest deviations are confined to the high-energy tail, where the number of contributing cells is small and the statistical fluctuations are correspondingly larger. The excess of energy could be due to the electromagnetic subsystem of the detector, since the ECal has a smaller cell size. This could affect the total energy distribution. However, the results of the total visible energy presented in figure 7 show that the effect of this overestimation is small.
The longitudinal shower profile (figure 4, center) describes how much energy is deposited on average in each of the 78 calorimeter layers. The first 30 layers are part of the ECal, while the last 48 compose the HCal. The model describes the distribution well. It can be observed that there are small bumps in the ECal layers, where the Geant4 simulation differs between even and odd layers. This alternating structure of higher and lower energy depositions per layer arises from the fact that for technical reasons, pairs of silicon sensors are placed on either side of one tungsten absorber layer facing opposite directions. These slabs are then installed into a tungsten structure containing every other absorber layer. The different amount of nonsensitive material results in the observed pair-wise difference in the mean energy between consecutive layers. While Geant4 reproduces this effect, CaloHadronic shows noticeable difficulties in modeling the subtle alternating pattern. This reflects a small limitation in capturing layer-dependent shower modulations. However these mismodelings only have a small effect on reconstruction for hadronic showers, as it is mostly a limitation for electromagnetic showers for which there are alternative generative models.
The radial shower profile (figure 4, right) describes the average radial energy distribution around the central shower axis (in the $y$-direction) in both ECal and HCal. CaloHadronic has a tendency to generate slightly narrower showers than Geant4. The shower profile is presented in appendix E in figure 20. Although these deviations are small, the slightly narrower showers produced by the model compared to Geant4 could reduce the separation power of particle flow object (PFO) reconstruction, particularly in environments where accurate discrimination between overlapping showers is critical. Nevertheless, CaloHadronic shows overall very good agreement with Geant4.
@@
In figure 7 the energy sum (total visible energy) for 5k showers at three incident energies for ECal, HCal and both calorimeters combined is shown. As noted earlier in the cell energy spectrum shown in Figure 4, CaloHadronic tends to slightly overestimate the hit energy, which could in principle affect the total energy distribution. However, the results here show that the impact of this overestimation is very minor.
@@ 4.2 Correlation Studies
Figure 11 shows two 2D histograms of the total visible energy deposited in the ECal against the total visible energy deposited in the HCal for incident energies of 15 GeV, 50 GeV and 85 GeV, respectively. The 1D histogram on the right side gives a clearer view of the HCal total visible energy when the ECal total visible energy is below 38 MeV. Most of the pion showers lie in the region of lower energy depositions in the ECal and higher energy depositions in the HCal. This is due to the fact that a large fraction of the pions travel through the ECal before showering in the HCal, where the majority of their energy is deposited. Compared to Geant4, CaloHadronic shows a slightly broader band of fluctuations and a slightly different slope in the visible energy distribution of ECal versus HCal. While this could, in principle, affect electron–hadron separation, in practice it does not pose an issue, as high-granularity calorimeters offer several alternative methods to achieve this discrimination.
@@ 5.1 Methods
In order to be able to apply the standard reconstruction suite of ILD to showers produced by the model, it is necessary to be able to correctly place the energy deposits into the detector geometry. We therefore use the DDML library [59, 60] to combine the model output with a full simulation application. We fire a $\pi^{+}$ with an energy of $50$ GeV from a particle gun positioned at (5,0,15-5,0,-15) mm [^6] This position is chosen to avoid the TPC cathode, which is positioned at $z=0$ in the global ILD coordinate system in the global ILD coordinate system, with a direction chosen to produce a perpendicular incidence at the calorimeter face. This takes into account the slight curvature of the track. The particles are not produced directly at the face of the calorimeter, as it is necessary for a track to be present in the tracking system for PandoraPFA to apply the correct sequence of algorithms for $\pi^{+}$ reconstruction. However, the energy of the pion used for conditioning of the model is extracted directly at the face of the calorimeter, to match the training scenario.
@@ 5.2 Results
6 Discussion
Overall, CaloHadronic reproduces the main features of hadronic showers with very good agreement to Geant4. The longitudinal and radial profiles, as well as the center of gravity distributions, all show that the model captures the essential shower development. Small deviations are visible, but they remain within acceptable limits for physics applications as demonstrated by the excellent post-reconstruction performance at the single shower level.
In the longitudinal profile, the alternating structure in the ECal layers is well described by Geant4 but slightly less so by CaloHadronic. While this effect is relevant for the reconstruction of electromagnetic showers, for which there are multiple other generative models, it only has a negligible effect on reconstruction for hadronic showers.
In the radial direction, CaloHadronic tends to generate showers that are slightly narrower than those from Geant4. This could, in principle, reduce the separation power of particle flow reconstruction where showers overlap.
Looking at the center of gravity (CoG), the model describes the distributions in all three directions ($x$, $z$, $y$) well. Minor differences are observed: the transverse CoGs ($x$, $z$) are slightly narrower, while along the shower axis ($y$) the distribution is shifted by a small amount.
As noted in the cell energy spectrum (figure 4), CaloHadronic slightly overestimates the hit energy, raising the concern that this might distort the total energy distribution. The comparisons presented later in figure 6 show that the impact of this effect is very subtle: while present, it does not significantly alter the global energy of the showers.
The number of hits and total visible energy per shower are also modeled with high fidelity across a wide range of incident energies. Both the resolution and linearity are preserved, and the comparison at fixed energies (15, 50, 85 GeV) confirms that CaloHadronic follows the Geant4 reference closely.
Correlation studies provide a more stringent test. The joint distributions of observables, such as energy sum versus center of gravity or number of hits versus energy sum, show overall agreement. The difference plots reveal systematic trends: CaloHadronic sometimes overpopulates the central regions while underpopulating the tails, and at fixed numbers of hits it slightly underestimates the energy sum compared to Geant4. Still, these deviations are modest, and the main correlations are well preserved.
The comparison of visible energy deposited in the ECal and HCal demonstrates that CaloHadronic captures the characteristic pion behavior. The model reproduces the global slope and fluctuations, with only slightly larger variations than Geant4. Importantly, while such differences could affect electron–hadron separation in principle, modern highly granular calorimeters provide several independent handles to achieve this task, so in practice it is not a limiting factor.
After downstream processing with PandoraPFA, CaloHadronic reproduces key reconstructed observables—including PFO energy, momentum, multiplicities, and particle ID—with accuracy comparable to Geant4. Both models show similar reconstruction inefficiencies, and at the event level CaloHadronic closely matches Geant4 for all PFOs as well as the leading PFO. This confirms that small deviations seen at the hit and shower level do not propagate into significant mismodeling after reconstruction at the single particle level. Future work could study the performance of CaloHadronic in multi-particle environments, where overlapping showers would provide more stringent tests of model performance. Moreover, incorporating the time dimension would be particularly relevant for studying hadronic shower development and pile-up effects. Ultimately, the model should be tested at the level of observables for higher level reconstructed objects (e.g. jets) in full physics events.
On CPU, CaloHadronic achieves a modest speed-up over Geant4, while on GPU the gain reaches up to three orders of magnitude. Higher numbers of function evaluations (NFE) reduce this advantage, but even with 59 NFE the GPU remains faster. Most of the runtime comes from the transformer-based diffusion modules, suggesting that accelerated sampling techniques could bring further improvements.
Taken together, these results show that CaloHadronic is able to reproduce both single-variable distributions and correlations with good accuracy. The small deviations observed are well understood and do not compromise the usefulness of the model for realistic physics studies. Importantly, CaloHadronic generates visually faithful shower shapes, particularly sub-shower tracks, which are crucial for Pandora Particle Flow to correctly reconstruct hadronic showers.
@@ 7 Conclusion
To turn generative models from proof-of-concept studies into useful surrogates to be used in production, they need to be able to handle the various complexities of realistic calorimeters. This work demonstrates for the first time how a point-cloud based generative model can be extended to simultaneously simulate the ECal and HCal parts of a hadronic shower, emphasizing the flexibility of this approach. Both individual shower properties, as well as correlations (also across ECal and HCal components), are described well. A classifier test shows that — as for the generation of partial showers — the generative fidelity needs to be further improved. This is especially visible in the fine grained track-like structures observed in the HCal. However the degree of agreement between generative output and ground-truth processing with reconstruction algorithms is of a similar or slightly better quality however the degree of agreement between generative output and ground-truth is of a similar or slightly better quality when also considering down-stream processing with reconstruction algorithms.
@@ 8 Acknowledgments
We thank William Korcari for early comments to this work. We want to thank Freya Blekman for valuable comments on the manuscript. This research was supported in part by the Maxwell computational resources operated at Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany. This project has received funding from the European Union’s Horizon 2020 Research and Innovation program under Grant Agreement No 101004761. We acknowledge support by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and via the KISS consortium (05D23GU4, 13D22CH5) funded by the German Federal Ministry of Research, Technology and Space (BMFTR) in the ErUM-Data action plan. P.M. has benefited from support by the CERN Strategic R&\&D Programme on Technologies for Future Experiments [45].
@@ Appendix E CaloHadronic further histograms
Figure 20: Distribution of hit $x$ (left), $y$ (middle), and $z$ (right) positions calculated from $50000$ showers. All distributions are calculated with a uniform distribution of incident particle energies between 10 and 90 GeV. The error band corresponds to the statistical uncertainty in each bin.